A mole of ideal polyatomic gas at 0°C and 1.00 atm is warmed up to expand isobarically to increase its volume by a factor of six. How much heat (in J) is transferred during the process?
Answer #1

Then , T1 = 273 K

V2/V1 = 6 and T2 = 6*273 = 1638K

Slufion> then T = 273K 3 - 6 Then from Charles Law, Tz = 6x273 = 1638 K Isobaric process the heat tranfered is, @=nCp dt =nc